• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Kumbartzki

    Log entry time 16:09:19 on July29,2006

    Entry number 176009

    keyword=Day shift summary

    Start with Radiator test on D target (15 cm)
    Run 1808 No radiator
    Run 1809 3%
    Run 1810 4%
    Run 1811 5%
    Next changed the beam position +- 1mm
    Operator has problems to get it right: first setting
    BPMA X 1.246 Y -0.532
    BPMB X -0.996 Y -0.017 moving around
    Run 1812
    Asked for beam 1 mm below the normal setting:
    BPMA X 1.269 Y -0.901
    BPMB X -1.005 Y -1.766
    Run 1813
    9:12 strart 1. optics run on C , centered the beam, Run 1814& 1815
    2. run 4 cm dummy
    3. run 15 cm dummy
    4. run Ta
    10:10 Ready for BPM calibration.
    Could do only the zero position. Had problems with the harp interface
    The IHA1H03B stopped short (see entry # 175964). Julie decided at this point not to continue with the BPM calibrations.
    11:30
    switch over to production run: set LD2, radiator 4%, raster on, 5uA
    run 1823 was stopped when we lost beam.
    Resume beam at 12:34, start run 1824. Beam currents 5 and 7uA,which gives 2.3KHz T3 and 25% deadtime for the left arm.
    Data runs 1825 - 1831
    15:25 Hall C makes access. We go for the BCM calibration
    Emily takes over.
    BCM is still in progress